OpenStudy (egbeach):

Find the limit of the function by using direct substitution. lim(2e^x cos x) x-> pi/2

OpenStudy (freckles):

do you have a problem pluggin in pi/2?

OpenStudy (anonymous):

can you two help me

OpenStudy (egbeach):

i just have no idea how to solve this problem

OpenStudy (freckles):

\[\lim_{x \rightarrow \frac{\pi}{2}}2e^x \cos(x)= 2 e^{\frac{\pi}{2}} \cos(\frac{\pi}{2})\] by direct substitution we can use direct substitution since the function exists at the thing the x is approaching

OpenStudy (anonymous):

im on the problem below

OpenStudy (freckles):

can you finishing simplifying

OpenStudy (egbeach):

thats the part i am having problems with.

OpenStudy (freckles):

so you aren't sure what cos(pi/2)=?

OpenStudy (egbeach):

-1/2

OpenStudy (freckles):

how do you get that

OpenStudy (freckles):

you can use the unit circle of the calculator if you want

OpenStudy (freckles):

cos(pi/2) should be zero either way

OpenStudy (egbeach):

oh

OpenStudy (freckles):

are you good?

OpenStudy (freckles):

\[\lim_{x \rightarrow \frac{\pi}{2}}2e^x \cos(x)= 2 e^{\frac{\pi}{2}} \cos(\frac{\pi}{2}) \\ =2e^{\frac{\pi}{2}} (0)=?\]